OpenStudy (anonymous):

x^2-13x+40=0 what are the solutions ?

OpenStudy (anonymous):

x=5 v x=8 ... typo :P simply use Quadratic Formula

OpenStudy (anonymous):

(x-8) (x-5) = 0 x=5 x=8 You should really try Wolfram alpha http://www.wolframalpha.com/input/?i=x^2-13x%2B40%3D0#

OpenStudy (anonymous):

Factor using which ever preferred method. Since is starts with x^2 each binomial will start with x. since the second +\- symbol is + the symbols will be the same and they will be the first +\- symbol (-) (x- )(x- ) now we need 2 numbers that multiply to give 40 and add to give us 13. Thats 5 and 8. (x-5)(x-8)=0 set each equal to 0 and solve. x-5=0, x=5 x-8=0. x=8 hope this helps